OpenStudy (izzy18970):

Can someone help me on this ??? Yolanda has 8 less hats than Garrett. Determine whether the equations accurately represent the relationship between y, the number of hats Yolanda has, and g, the number of hats Garrett has. Expression y=g+8 y=g−8 g=y+8 g=y−8 Part 2 Yolanda has 13 hats. How many hats does Garrett have?

OpenStudy (izzy18970):

yes or no

OpenStudy (izzy18970):

for the expressions

OpenStudy (anonymous):

These two are correct: y=g−8 g=y+8 Part 2: substitute 13 in for y g=13+8 g=21 garret has 21 hats!

OpenStudy (izzy18970):

thx